# Jiazhen Liu, Peihan Li, Yuwei Wu, Gaurav S. Sukhatme, Vijay Kumar, Lifeng Zhou

### v2

@@ 3.3 Reformulated Optimization
$~{}~{}~{}~{}~{}\quad g_{\mathcal{S}_{l}}(\mathbf{x}_{i})\geq 0,\forall i\in[M],\forall l\in[p],$ (22c)
$~{}~{}~{}~{}~{}\quad h_{\mathcal{C}_{k}}(\mathbf{x}_{i})\geq 0,\forall i\in[M],\forall k\in[q].$ (22d)
We use Forces Pro [22, 23] to solve this nonlinear optimization and obtain control inputs for the robots at every time step.
@@ 4 Simulations
We evaluate the performance of our risk-aware target tracking framework under the setting where (i) there exists a sensing danger zone; and (ii) there exists a communication danger zone in MATLAB simulations111 We open source our code at: https://github.com/Zhourobotics/DZone_Tracking.
@@ 4.1 Sensing danger zones
The qualitative results containing trajectories of both robots and targets as shown in Fig. 2. Different condition parameters impact robots’ trajectories when they are in proximity to the sensing danger zone. The subfigures in the first and second rows illustrate that the robots’ behaviors become more conservative to keep a larger distance from the danger source with the increase in the uncertainty $\mathbf{\Sigma}$. Compared with the first parameter setting (first row), the third one decreases the value of $\epsilon_{1}$, i.e., it imposes stricter risk requirements on robots prohibiting them from entering the sensing danger zone. In this case, it shows that the robots also keep a larger distance away from the danger source.
@@
(a) Trace
(b) Probability
Figure 3: The trace (a) and the probability of sensor failure (b) throughout the tracking process using three different sets of parameters ($\mathbf{\Sigma}$, $\epsilon_{1}$).
@@ 4.2 Communication danger zones
We also evaluate the tracking performance under one communication danger zone to illustrate how an uncertain communication danger zone influences robots’ risk-aware behavior. In a setting where 4 robots track 2 targets, we adopt three different combinations of parameters including the uncertainty of the jamming source’s position $\mathbf{\Sigma}$ and risk requirement $\epsilon_{2}$ to compare robot trajectories. The two targets have a circular motion inside a communication danger zone in a counter-clockwise direction. The weights in the objective function (refer to Eq. 8a) are $w_{1}=2.0,w_{2}=0.002$.
@@
The corresponding quantitative results, consisting of trace and the probability of communication jamming, are presented in Fig. 5. Fig. 5(a) verifies that either increasing the uncertainty level $\mathbf{\Sigma}$ or decreasing the required risk level $\epsilon_{2}$ steer robots away from the communication danger zone. This prevents them from closely following the targets, and consequently, decreases the tracking performance. The probability of jamming as shown in Fig. 5(b) is estimated through sampling. Specifically, we sample the position of the jamming source 1000 times from the Gaussian distribution it follows. For a particular robot $i$, we calculate its distance to the sampled position, $a$, and its distance from its furthest teammate $c^{*}$ (see Eq. 6). Particularly, we use only teammates that are within a communication range, i.e., neighbors, to calculate this distance. Then, for each sample, we test whether $a<\delta_{2}c^{*}$ holds. If the condition holds, we consider robot $i$ to be jammed. Otherwise, we consider it to be safe. Supposing the condition check holds for $\beta$ times, then the estimated probability is $\frac{\beta}{1000}$. We average the probability value across all robots to obtain the results plotted in Fig. 5(b). As seen, the approximated probability of communication jamming remains below the required upper-bound $\epsilon_{2}$ (shown in dotted lines) for all three cases. Notably, for the third case, the approximated probability remains nearly 00 throughout the whole tracking process since the risk requirement $\epsilon_{2}=0.02$ is stringent.
(a) Trace
(b) Probability
Figure 5: The trace (a) and the probability of communication jamming (b) throughout the tracking process using three different sets of parameter $(\mathbf{\Sigma},\epsilon_{2})$.
@@ 5 Experiments
We conducted several hardware experiments to evaluate the performance under real-world tracking uncertainty and danger environments. We use a team of Crazyflie drones [24] as trackers and multiple Yahboom ROSMASTER X3 ground robots [25] as targets. Danger zones are set up as paper disks on the floor of an indoor environment. We consider the ground targets with non-adversarial movements. The target trajectories are prespecified but the trackers are agnostic to the movement pattern. Trackers need to estimate the positions of ground targets and follow them closely while satisfying the risk requirements. The optimal control inputs are solved online in Eq. 22 at every time step on a local Desktop and sent to Crazyflies. Eq. 22 is solved within 50ms on average. Since we are not imposing inter-robot collision avoidance constraints, we make the trackers fly at different fixed heights.
We validated the scenarios in simulations from Sec. 4 on the real-robot system and show the risk-aware tracking behaviors in Fig. 6. The parameters including $\epsilon_{1}$, $\epsilon_{2}$, $\mathbf{\Sigma}$, and the safety clearance of the sensing danger zone are adjusted to suit the indoor experimental conditions. Fig. 6(a) and Fig. 6(b) display the robots’ behaviors around a sensing and a communication danger zone, respectively. The robots demonstrate risk awareness in their actions. When the targets are inside the sensing danger zone, robots take a detour and keep a distance to ensure their safety. Similarly, when targets approach the communication danger zone, pairs of robots tracking the same target stay close to each other to secure their communication links.
@@
During the experiments, due to imprecise localization and actuation uncertainty, we occasionally run into situations where trackers end up being within a danger zone at some time steps and the reformulated optimization in Eq. 22 become infeasible. This demonstrates the difficulty of sim-to-real transfer for multi-robot target tracking. To address this situation, we apply a single-step control input such that the trackers can resiliently escape from danger zones. The escape control points towards the direction from the mean position of the danger source to tracker’s current position. After escaping from the zone, trackers resume the regular planning procedure.
@@ 6 Conclusion and Discussion
In this paper, we proposed a risk-aware multi-robot target tracking framework in danger zones. We modeled the danger zones into different attack types and formulated the optimal control as a nonlinear optimization problem. The problem is reformulated with approximated chance constraints and tracking mission objectives. We verified in both simulation and hardware experiments that with different uncertainty levels and risk requirements, robots exhibit corresponding risk-aware behaviors in proximity to danger zones. For real-world deployments with the condition that robots fall into the danger zones, we provide a practical method to escape the danger zone and continue the tracking task. In the future, we will extend the works for resilient tracking and address the perception and joint localization in target tracking [26] for better sim-to-real transfer.

### v3

@@ 3.3 Reformulated Optimization
$~{}~{}~{}~{}~{}\quad g_{\mathcal{S}_{l}}(\mathbf{x}_{i,t+1})\geq 0,\forall i\in[M],\forall l\in[p],$ (19c)
$~{}~{}~{}~{}~{}\quad h_{\mathcal{C}_{k}}(\mathbf{x}_{i,t+1})\geq 0,\forall i\in[M],\forall k\in[q].$ (19d)
We use a nonlinear solver Forces Pro [22, 23] to solve this nonlinear optimization and obtain control inputs for the robots at every time step.
@@ 4 Simulations
We evaluate the performance of our risk-aware target tracking framework under the setting where (i) there exists a sensing danger zone; (ii) there exists a communication danger zone, in MATLAB simulations111 We open source our code at: https://github.com/Zhourobotics/DZone_Tracking.
@@ 4.1 Sensing danger zones
The qualitative results containing trajectories of both robots and targets are shown in Fig. 2. Different condition parameters impact robots’ trajectories when they are in proximity to the sensing danger zone. The subfigures in the first and second rows illustrate that the robots’ behaviors become more conservative to keep a larger distance from the danger source with the increase in uncertainty $\mathbf{\Sigma}$. Compared with the first parameter setting (first row), the third one decreases the value of $\epsilon_{1}$, i.e., it imposes stricter risk requirements on robots prohibiting them from entering the sensing danger zone. In this case, it shows that the robots also keep a larger distance away from the danger source.
@@
(a) Trace (S)
(b) Probability (S)
(c) Trace (C)
(d) Probability (C)
Figure 3: (a)(b): trace and probability of sensor failure throughout the tracking process with a Sensing danger zone; (c)(d) trace and probability of communication jamming with a Communication danger zone. The comparison of results under three sets of parameters is shown.
@@ 4.2 Communication danger zones
We also evaluate the tracking performance under one communication danger zone to illustrate how an uncertain communication danger zone influences robots’ risk-aware behavior. In a setting where 4 robots track 2 targets, we adopt three different combinations of parameters including the uncertainty of the jamming source’s position $\mathbf{\Sigma}$ and risk requirement $\epsilon_{2}$ to compare robot trajectories. The two targets follow a circular motion inside a communication danger zone. The weights in the objective function (refer to Eq. 8a) are $w_{1}=2.0,w_{2}=0.002$.
@@
The corresponding quantitative results, consisting of trace and the probability of communication jamming, are presented in Fig. 3. Fig. 3(c) verifies that either increasing the uncertainty level $\mathbf{\Sigma}$ or decreasing the required risk level $\epsilon_{2}$ steer robots away from the communication danger zone. This prevents them from closely following the targets, and consequently, decreases the tracking performance. The probability of jamming as shown in Fig. 3(d) is estimated through sampling. Specifically, we sample the position of the jamming source 1000 times from the Gaussian distribution it follows. For a particular robot $i$, we calculate its distance to the sampled position, $a$, and its distance from its furthest teammate $c^{*}$ (see Eq. 6). Particularly, we use only teammates that are within a communication range, i.e., neighbors, to calculate this distance. Then, for each sample, we test whether $a<\delta_{2}c^{*}$ holds. If the condition holds, we consider robot $i$ to be jammed. Otherwise, we consider it to be safe. Supposing the condition check holds for $\beta$ times, then the estimated probability is $\frac{\beta}{1000}$. We average the probability value across all robots to obtain the results plotted in Fig. 3(d). As seen, the approximated probability of communication jamming remains below the required upper-bound $\epsilon_{2}$ (shown in dotted lines) for all three cases. Notably, for the third case, the approximated probability remains nearly 00 throughout the whole tracking process since the risk requirement $\epsilon_{2}=0.02$ is stringent.
@@ 5 Experiments
We conducted several hardware experiments to evaluate the performance under real-world tracking uncertainty and dangerous environments. We use a team of Crazyflie drones [24] as trackers and multiple Yahboom ROSMASTER X3 ground robots [25] as targets. Danger zones are set up as paper disks on the floor of an indoor environment. We consider the ground targets with non-adversarial movements. The target trajectories are prespecified but trackers are agnostic to the movement pattern. Trackers need to estimate the positions of ground targets and follow them closely while satisfying the risk requirements. The optimal control inputs are solved online in Eq. 19 at every time step on a local Desktop and sent to Crazyflies. Eq. 19 is solved within 50ms on average. Since we are not imposing inter-robot collision avoidance constraints, we make the trackers fly at different fixed heights.
We validated the simulation scenarios from Sec. 4 on the real-robot system and showed the risk-aware tracking behaviors in Fig. 5. The parameters including $\epsilon_{1}$, $\epsilon_{2}$, $\mathbf{\Sigma}$, and the safety clearance of the sensing danger zones are adjusted to suit the indoor experimental conditions. Fig. 5(a) and Fig. 5(b) display the robots’ behaviors around a sensing and a communication danger zone, respectively. The robots demonstrate risk awareness in their actions. When the targets are inside the sensing danger zone, robots take a detour and keep a distance to ensure their safety. Similarly, when targets approach the communication danger zone, pairs of robots tracking the same target stay close to each other to secure their communication links.
@@
During the experiments, due to imprecise localization and actuation uncertainty, we occasionally run into situations where trackers end up being within a danger zone at some time steps and the reformulated optimization in Eq. 19 become infeasible. This demonstrates the difficulty of sim-to-real transfer for multi-robot target tracking. To address this situation, we apply a single-step control input such that the trackers can resiliently escape from danger zones. The escape control points towards the direction from the mean position of the danger source to tracker’s current position. After escaping from the zone, trackers resume their regular planning procedure.
@@ 6 Conclusion and Discussion
In this paper, we proposed a risk-aware multi-robot target tracking framework in danger zones. We modeled the danger zones into different attack types and formulated the optimal control as a nonlinear optimization problem. The problem is reformulated with approximated chance constraints and tracking mission objectives. We verified in both simulations and hardware experiments that with different uncertainty levels and risk requirements, robots exhibit corresponding risk-aware behaviors in proximity to danger zones. For real-world deployments with the condition that robots fall into the danger zones, we provide a practical method to escape from the danger zone to continue tracking task. In the future, we will extend the work for resilient tracking and address the perception and joint localization in target tracking [26] for better sim-to-real transfer.